OpenStudy (anonymous):

The perimeter of a rectangle is 54 cm. The area of the same rectangle is 176 cm². What are the dimensions of the rectangle? a. 11 cm by 16 cm b. 8 cm by 22 cm c. 5.5 cm by 332 cm d. 4 cm by 44 cm

OpenStudy (anonymous):

11 cm by 16 cm. i think..

OpenStudy (anonymous):

2x+2y=54 x*y =176 x(27-x) = 176 -x^2+27x = 176 -(x-13.5)^2+182.25=176 (x-13.5)^2 = 6.25 x-13.5 = ±2.5 x = 16 or 11 y = 176/16 = 11